a chemist needs 13 liters of a 10% acid solution. How much of a 5% solution should be mixed with a 31% solution to get that?
Answer by richwmiller(17219)

a+b=13,
0.05*a+0.31*b=0.1*13
a=13-b
0.05*(13-b)+0.31*b=1.3
0.65-0.05b+0.31*b=1.3
0.26*b=0.65
b=2.5
a=13-b
a=10.5 liter at 5%
b=2.5 liter at 31%
check
0.05*10.5+0.31*2.5=0.1*13
0.525+0.775=1.3
1.3=1.3
ok